OpenStudy (anonymous):

what is the differentiation rule for (fg)(1)

OpenStudy (anonymous):

sorry i didnt follow

OpenStudy (anonymous):

well (fg)'(x) is f'(x) x g(x) + f(x) x g'(x) when its written like (fg)(x)...what do you do

OpenStudy (anonymous):

same as before

OpenStudy (anonymous):

satellite that didnt work for me

OpenStudy (anonymous):

dont know why

OpenStudy (anonymous):

maybe you can post the exact question

OpenStudy (anonymous):

fg(x)=(fog)(x)=f'[g(x)]*g'(x) so fg(1)= f'[g(1)]*g'(1

OpenStudy (anonymous):

ok let met just do that here

OpenStudy (anonymous):

nvm i think barbillus got it right

OpenStudy (anonymous):

aaah \[f\circ g\] got it!

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Differentiation_rules

OpenStudy (anonymous):

barbillus if at at 1 i have f(x)= 1 f'(x)=-1 g(x)=2 g'(x)=3

OpenStudy (anonymous):

am i doing -1(2) x (3)

OpenStudy (anonymous):

f'(2)*3=-1*3=-3

OpenStudy (anonymous):

my software rejected that

OpenStudy (anonymous):

f'(g(x))*g'(x) ; f'(g(1))*g'(1) = f'(2)*3=-1*3=3 should be -3

OpenStudy (anonymous):

pls look at that

OpenStudy (anonymous):

ok let me check

OpenStudy (anonymous):

it asks for fg(1) not its derivative f'g(1) so we should put f(g(1))=f(2)=-3 again -3 i dunno why it is incorrect

OpenStudy (anonymous):

i came up with -3 earlier and i agree with you

OpenStudy (anonymous):

oh well thanks for the help

OpenStudy (anonymous):

yea it's definitely -3

OpenStudy (anonymous):

ill post it again on the main forum and see if anyone comes up with anything different

OpenStudy (anonymous):

let me know if you get sth pls

OpenStudy (anonymous):

what exactly is sth

OpenStudy (anonymous):

or what does it mean

OpenStudy (anonymous):

i mean please inform me

OpenStudy (anonymous):

ok. im going to post it on the main forum. you just click and watch and see if anyone find anything different

OpenStudy (anonymous):

or i can write here if someone get anything different

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

pipa go here http://openstudy.com/study#/updates/4f3b2dade4b0fc0c1a0ef37a

OpenStudy (anonymous):

barbillus, the rule for (fg)(x) is just f x g. so the correct answer is 2 http://www.wolframalpha.com/input/?i=differentiate+%28fg%29%28x%29

OpenStudy (anonymous):

oh thanks, i knew it as (fog)(x),,yeah then the result is 2